OpenStudy (anonymous):

You deposit 6000 In an account that pays 5.25% interest compounded semiannually. After 10 years, the interest rate is increased to 5.4% compounded quarterly. What will be the value of the account after 18 years?

OpenStudy (anonymous):

is this a situation where you'd use I = prt ?

OpenStudy (anonymous):

Use the formula A=P(1+ r/n)^(nt) P=principal r= rate of interest as a decimal n= number of times compounded in a year t= number of years

OpenStudy (anonymous):

multipy 6000 * 5.25+5.4 * 18

OpenStudy (anonymous):

5.25+5.4=10.65 6000* 10.65 * 18 and you'll get your answer now can I have a fan or a medal?

OpenStudy (anonymous):

@Starr_DynastyT help openstudy might be down because I can't get anyone's questions so I can't help anyone and it's killing me